# Toward Reliable Biomedical Hypothesis Generation: Evaluating Truthfulness and Hallucination in Large Language Models

### v1

@@ Abstract
Large language models (LLMs) have shown significant potential in scientific disciplines such as biomedicine, particularly in hypothesis generation, where they can analyze vast literature, identify patterns, and suggest research directions. However, a key challenge lies in evaluating the truthfulness of generated hypotheses, as verifying their accuracy often requires substantial time and resources. Additionally, the hallucination problem in LLMs can lead to the generation of hypotheses that appear plausible but are ultimately incorrect, undermining their reliability. To facilitate the systematic study of these challenges, we introduce TruthHypo, a benchmark for assessing the capabilities of LLMs in generating truthful biomedical hypotheses, and KnowHD, a knowledge-based hallucination detector to evaluate how well hypotheses are grounded in existing knowledge. Our results show that LLMs struggle to generate truthful hypotheses. By analyzing hallucinations in reasoning steps, we demonstrate that the groundedness scores provided by KnowHD serve as an effective metric for filtering truthful hypotheses from the diverse outputs of LLMs. Human evaluations further validate the utility of KnowHD in identifying truthful hypotheses and accelerating scientific discovery. Our data and source code are available at https://github.com/Teddy-XiongGZ/TruthHypo.
@@ 1 Introduction
One particularly promising application of LLMs is their use in biomedical hypothesis generation, where they can assist in identifying promising research directions (Park et al., 2024; Si et al., 2024; Guo et al., 2025). By analyzing extensive biomedical literature, LLMs can uncover gaps in existing knowledge and propose novel hypotheses that may not be immediately apparent to human researchers. For instance, LLMs have been successfully applied to propose novel drug combinations for breast cancer treatment, some of which were later validated in laboratory experiments, showcasing their potential to accelerate biomedical discoveries (Abdel-Rehim et al., 2024).
Despite these advancements, there are substantial challenges that limit the practical utility of LLMs in biomedical hypothesis generation. A critical concern is the inability to evaluate the truthfulness of generated hypotheses. While LLMs can generate hypotheses that seem plausible, it remains uncertain whether these hypotheses are valid and grounded in existing knowledge or merely hallucinated and scientifically invalid. This issue is further exacerbated by the well-documented “hallucination” problem, where LLMs confidently produce information that is factually inaccurate or unsupported, posing challenges to their reliability in biomedical contexts (Jin et al., 2024). While current research has largely focused on improving the novelty and diversity of LLM-generated hypotheses, their truthfulness and grounding in established knowledge remain underexplored (Baek et al., 2024; Hu et al., 2024; Si et al., 2024).
@@
To address these challenges, we introduce TruthHypo, a comprehensive benchmark for evaluating the ability of LLMs to generate truthful biomedical hypotheses, and KnowHD, a knowledge-based hallucination detection framework designed to assess the groundedness of these hypotheses. TruthHypo, built on a biomedical knowledge graph along with a domain-specific corpus, provides a controlled environment to evaluate how well LLM-generated hypotheses align with established biomedical knowledge. KnowHD focuses on analyzing the reasoning processes of LLMs to identify hypotheses that are likely hallucinated or untruthful. Our findings reveal that LLMs face significant challenges in generating truthful hypotheses. By analyzing hallucinations in the reasoning processes behind generated hypotheses, we demonstrate that groundedness scores from KnowHD serve as an effective signal for identifying truthful hypotheses from the diverse outputs of LLMs. Human evaluations on open-ended hypothesis generation tasks further confirm the utility of KnowHD in identifying scientifically valid hypotheses.
@@
  • We introduce TruthHypo, a comprehensive benchmark designed to evaluate the ability of LLMs to generate truthful biomedical hypotheses.
@@ 7 Conclusion
We presented TruthHypo, a benchmark for evaluating the ability of LLMs to generate truthful biomedical hypotheses, and KnowHD, a framework for detecting hallucinations by assessing groundedness in reasoning. Through extensive evaluation, we highlighted the limitations of existing LLMs and demonstrated that selecting highly grounded hypotheses improves truthfulness. These contributions offer valuable insights for improving the reliability and utility of LLMs in scientific discovery.

### v2

# Toward Reliable Scientific Hypothesis Generation: Evaluating Truthfulness and Hallucination in Large Language Models

@@ Abstract
Large language models (LLMs) have shown significant potential in scientific disciplines such as biomedicine, particularly in hypothesis generation, where they can analyze vast literature, identify patterns, and suggest research directions. However, a key challenge lies in evaluating the truthfulness of generated hypotheses, as verifying their accuracy often requires substantial time and resources. Additionally, the hallucination problem in LLMs can lead to the generation of hypotheses that appear plausible but are ultimately incorrect, undermining their reliability. To facilitate the systematic study of these challenges, we introduce TruthHypo, a benchmark for assessing the capabilities of LLMs in generating truthful scientific hypotheses, and KnowHD, a knowledge-based hallucination detector to evaluate how well hypotheses are grounded in existing knowledge. Our results show that LLMs struggle to generate truthful hypotheses. By analyzing hallucinations in reasoning steps, we demonstrate that the groundedness scores provided by KnowHD serve as an effective metric for filtering truthful hypotheses from the diverse outputs of LLMs. Human evaluations further validate the utility of KnowHD in identifying truthful hypotheses and accelerating scientific discovery. Our data and source code are available at https://github.com/Teddy-XiongGZ/TruthHypo.
@@ 1 Introduction
One particularly promising application of LLMs is their use in scientific hypothesis generation, where they can assist in identifying promising research directions (Park et al., 2024; Si et al., 2024; Guo et al., 2025). By analyzing extensive scientific literature, LLMs can uncover gaps in existing knowledge and propose novel hypotheses that may not be immediately apparent to human researchers. For instance, LLMs have been successfully applied to propose novel drug combinations for breast cancer treatment, some of which were later validated in laboratory experiments, showcasing their potential to accelerate biomedical discoveries (Abdel-Rehim et al., 2024).
Despite these advancements, there are substantial challenges that limit the practical utility of LLMs in scientific hypothesis generation. A critical concern is the inability to evaluate the truthfulness of generated hypotheses. While LLMs can generate hypotheses that seem plausible, it remains uncertain whether these hypotheses are valid and grounded in existing knowledge or merely hallucinated and scientifically invalid. This issue is further exacerbated by the well-documented “hallucination” problem, where LLMs confidently produce information that is factually inaccurate or unsupported, posing challenges to their reliability in scientific contexts (Jin et al., 2024). While current research has largely focused on improving the novelty and diversity of LLM-generated hypotheses, their truthfulness and grounding in established knowledge remain underexplored (Baek et al., 2024; Hu et al., 2024; Si et al., 2024).
@@
To address these challenges, we introduce TruthHypo, a comprehensive benchmark for evaluating the ability of LLMs to generate truthful scientific hypotheses, and KnowHD, a knowledge-based hallucination detection framework designed to assess the groundedness of these hypotheses. TruthHypo, built on a biomedical knowledge graph along with a domain-specific corpus, provides a controlled environment to evaluate how well LLM-generated hypotheses align with established scientific knowledge. KnowHD focuses on analyzing the reasoning processes of LLMs to identify hypotheses that are likely hallucinated or untruthful. Our findings reveal that LLMs face significant challenges in generating truthful hypotheses. By analyzing hallucinations in the reasoning processes behind generated hypotheses, we demonstrate that groundedness scores from KnowHD serve as an effective signal for identifying truthful hypotheses from the diverse outputs of LLMs. Human evaluations on open-ended hypothesis generation tasks further confirm the utility of KnowHD in identifying scientifically valid hypotheses.
@@
  • We introduce TruthHypo, a comprehensive benchmark designed to evaluate the ability of LLMs to generate truthful scientific hypotheses.
@@ 7 Conclusion
We presented TruthHypo, a benchmark for evaluating the ability of LLMs to generate truthful scientific hypotheses, and KnowHD, a framework for detecting hallucinations by assessing groundedness in reasoning. Through extensive evaluation, we highlighted the limitations of existing LLMs and demonstrated that selecting highly grounded hypotheses improves truthfulness. These contributions offer valuable insights for improving the reliability and utility of LLMs in scientific discovery.